Exercise 6-5 Companywide and Segment Break-Even Analysis [LO6-5] Piedmont Company segments its business into two regions-North and South. The company prepared the contribution format segmented income statement shown below: Total CompanyNorth South $800,000 $600,000 $200,000 560,000 480,000 80,000 Sales Variable expenses Contribution margin 240,000120,000 120,000 61,00061,000 Traceable fixed expenses Segment margin Common fixed expenses Net operating income 122,000 18,000 $ 59,000 $ 59,000 2,000 $ 66,000 Required: 1. Compute the companywide break-even point in dollar sales Dollar sales for company to break
media%2Feb8%2Feb8b8302-986c-443c-91e9-31
Answer #1

Contribution margin ratio = contribution /sales

North :120000/600000 = .20 or 20%

south :120000/200000= .60 or 60%

weight of north : 600000/800000= .75 or 75%              [sales of north /total sales]

weight of south : 200000/800000= .25 or 25%

weighted average contribution margin ratio : [20*.75]+[60*.25]

             = 15+ 15

                = 30%

Total fixed cost : 122000+52000= 174000

a)BEP for a company as a whole : fixed cost /weighted average contribution margin ratio

                            = 174000/ .30

                           = $ 580000

2)BEP(north ) = segment fixed cost /contribution margin ratio

         = 61000/.20

       = $ 305000

3)BEP (south) =61000 /.60

         = $ 101667